# Uncovering Hidden Connections: Iterative Search and Reasoning for Video-grounded Dialog

### v3

@@ Abstract
1 Introduction
@@
With the rapid proliferation of multimedia data, research attention has focused on vision-language tasks, proving their pivotal role in academic research [1, 2, 3, 4, 5, 6]. Notable advancements have been realized in fields such as Visual Question Answering (VQA) [7, 8, 9], Visual Dialog [10, 11], and Video Captioning [12, 13, 14]. Central to these tasks is the aim to refine machine capabilities in understanding visual content and to convey these understandings linguistically in ways intelligible to humans.
@@
An emerging domain of significance is video-grounded dialog [^1] https://video-dialog.com/, sometimes referred to as audio-visual scene-aware dialog by certain academicians [15, 16, 17, 18]. As illustrated in Fig. 1, this rapidly evolving domain mandates the creation of an agent adept at conducting coherent dialogs rooted in video data, which is a monumental leap toward genuine human-computer synergy. The implications of video-grounded dialogue extend widely, enhancing virtual personal assistants and creating innovative solutions for visually impaired individuals, highlighting its significant academic relevance. Spurred by advancements in deep learning, the domain of video-grounded dialog is witnessing considerable traction. Existing techniques in this domain can be methodically categorized into:
@@
1) Textual Information Modeling. This branch emphasizes modeling textual elements, primarily involving question inquiries and their preceding contexts [19, 20, 21]. For instance, Le et al. [20] devised a semantic graph based on the lexical components of question-answer pairs, leading to a model proficient in forecasting reasoning pathways on the graph. This method is adept at extracting question-specific data from dialog contexts. Similarly, Kim et al. [19] employed the Gumbel-Softmax technique to identify dialog content that is most pertinent to a given question, enhancing question representation in the process.
@@
2) Visual Information Modeling. This kind of method highlights visual elements, primarily focusing on video frames and notable objects [22, 23, 24, 25]. As a case in point, Le et al. [23] propounded a bidirectional model that encapsulates both spatial-to-temporal relations and their converse, aiming to mine the complex spatial and temporal cues embedded in the video. Concurrently, Geng et al. [22] advocated representing videos via spatio-temporal scene graphs, capturing salient audio-visual cues and their semantic interrelations effectively.
@@
3) Decoder Generation. This section emphasizes the formulation of answer derivations [26, 27, 28]. For example, Le et al. [27] extended the established functionalities of the GPT-2 architecture, discarding conventional auto-regressive decoding strategy. Complementing this domain, the integration of pointer networks [26] has amplified the generative capabilities of dialog systems.
@@
Despite the significant progress observed in the domain of video-grounded dialog, there are still critical challenges that demand meticulous consideration: 1) Dialog History Understanding. Notable recent studies [21, 20] predominantly focus on dialog history segments that align directly with user queries. Such a perspective can inadvertently exclude broader dialog contexts, leading to the omission of significant information. The inherent lack of transparency in these methods further impedes a clear understanding of the decision-making process of the model. Hence, the imperative lies in developing a superior textual encoder capable of thoroughly analyzing dialog histories while improving model interpretability. 2) Visual Content Understanding. While several works [22] endeavor to extract visual signals through complex graph structures, the importance of user-posed questions in steering visual comprehension is frequently disregarded. This omission leads to a failure to capture question-specific insights, thereby limiting the capacity to mine meaningful visual information. Additionally, the current trend of depending solely on a singular extraction process [24, 25] may be inadequate considering the multi-faceted nature of video data. Therefore, an approach that harmonizes user queries with iterative reasoning is crucial for a nuanced comprehension of visual content.
Figure 1: An illustration of video-grounded dialog.
@@
2 Related Work
2.1 Visual Dialog
@@
Visual dialog represents a sophisticated extension of the VQA framework, requiring multiple rounds of sequential interactions. The fundamental goal of this paradigm is to facilitate a seamless integration of cross-modal information derived from various inputs, culminating in the formation of a coherent textual response [29, 30]. Current methods within this domain may be broadly grouped into three distinct paradigms: attention mechanism-based strategies, graph network-based techniques, and approaches reliant on pre-trained models.
@@
Attention Mechanism-based Approaches. These approaches deploy attention mechanisms to pinpoint the visual regions that bear the highest relevance to the posed question. The underlying process requires an acute understanding of the immediate context of the inquiry, as well as an insight into the historical dialog trajectory. A salient example of this approach is the Recursive visual Attention (RvA) model, formulated by Niu et al. [31]. Through iterative refinement of visual attention and retrospective analysis of dialog sequences, this model adeptly addresses co-reference ambiguities. Further complementing this area, Yang et al. [32] have innovated two Sequential Dialog Networks (SeqDialN). These networks artfully separate the complexity inherent in multimodal feature fusion from the inference process, thereby simplifying the design and implementation of the inference engine.
@@
Graph Network-based Techniques. Recognizing the pivotal role of seamless information integration, the academic community is increasingly embracing graph network techniques [33]. For instance, Schwartz et al. [34] brought forth the Factor Graph Attention (FGA) architecture. This methodology enables the meticulous construction of graphs atop utility representations, subsequently discerning their interconnections. In a parallel vein, Guo et al. [35] introduced the Context-Aware Graph (CAG) neural structure, characterized by the dynamic interdependencies exhibited by individual nodes within the graph. Uniquely, this paradigm ensures that only nodes bearing the utmost relevance are integrated, yielding a refined, context-sensitive relational graph logic.
@@
Pre-trained Model-dependent Methods. The contemporary era of machine learning has been marked by the rise of pre-trained models, praised for their ability to seamlessly integrate diverse informational modalities. Anchoring on this insight, Wang et al. [36] delineated a holistic strategy that harnesses pre-trained architectures to adroitly weave together visual and dialogual elements. Mirroring this trajectory, Nguyen et al. [37] unveiled the Light-weight Transformer for Many Inputs (LTMI), a pioneering neural framework harnessing an agile transformer mechanism to masterfully orchestrate interactions across diverse utilities.
@@
However, the nuances and complexities of video-grounded dialog task preclude the direct application of methodologies cultivated for image-based visual dialog. Video-grounded dialog manifests longer utterance sentences, additional dialog rounds, and a more intricate context that demands sophisticated comprehension [38]. These attributes compound the challenge, rendering the direct transposition of existing visual dialog techniques impractical and inefficacious.
2.2 Video-grounded Dialog
@@
In recent years, the academic community has observed an escalating focus on video-grounded dialog systems. Noteworthy challenges, including DSTC7 [38] and DSTC8 [39], have been instrumental in formulating benchmarks in this domain. This burgeoning field has subsequently witnessed the advent of numerous sophisticated methodologies to address its inherent complexities. For example, Hori et al. [15] pioneered an approach employing an LSTM-driven encoder-decoder framework, emphasizing multimodal attention to harmoniously integrate textual and visual cues. Expanding on this work, Le et al. [16] presented the Multimodal Transformer Network (MTN), devised specifically to encode visual content while concurrently integrating diverse data modalities. Similarly, innovative contributions like the Bi-directional Spatio-Temporal reasoning model (BiST)[23] have made strides in extracting both visual and spatial cues from videos, thus elucidating the interconnectedness of text and visual sequences in multidimensional aspects. In pursuit of refining co-referential understanding, Kim et al. [19] introduced the Structured Co-reference Graph Attention model (SCGA), laying the foundation for a graph-based approach deeply rooted in multimodal co-referencing techniques. In tandem with this, Geng et al. [22] advanced the Spatio-Temporal Scene Graph Representation learning technique (STSGR)—an innovative proposition capturing the semantic crux of videos through scene graphs, all the while maintaining efficient encoding.
@@
The ascendance of pre-trained language models in myriad natural language processing domains is unmistakable. Recent efforts are directed toward fusing these models with video-grounded dialog systems. To this end, Le et al. [27] advanced the VGD-GPT framework, leveraging the capabilities of the well-regarded GPT-2 model, thus transforming video-grounded dialog generation into a more fluid sequence-to-sequence task. Building on this progression, Li et al. [17] incorporated a multi-task learning approach with pre-existing language models, culminating in the inception of the Response Language Model (RLM). Recognizing the hurdles of text hallucination in text generation, Yoon et al. [28] crafted the Text HallucinAtion Mitigating (THAM) architecture, integrating Text Hallucination Regularization (THR) loss into expansive pre-trained language models. In a stride towards achieving nuanced dialog generation, Zhao et al. [40] advocated for a Multi-Agent Reinforcement Learning (MARL) methodology, accentuating sophisticated collaborations in cross-modal analysis. With the advent of Large Video-Language Models (LVLMs), such as Video-LLaMA [41] and Otter [42], it is evident that this intricate domain continues to be at the forefront of research attention.
@@
3 Methodology
As illustrated in Figure 2, our model consists of three components: a textual encoder, a visual encoder, and a generator. The textual encoder is designed to perform interactions between questions and history contexts and output enhanced representations. Meanwhile, the visual encoder is utilized to implement progressive reasoning between questions and videos, enhancing video content comprehension. Finally, the generator combines multiple input representations to produce system answers. In the following sections, we will elaborate on each component in detail.
3.1 Problem Formulation
@@
3.2 Textual Encoder
3.2.1 Preprocessing
The word embeddings for the given question $Q_{t}$, the caption $C$, and the dialog history $H$ are obtained through the same process. For clarify, we consider the question $Q_{t}=(w_{1},w_{2},...,w_{n_{q}})$ as an example. Here, an operation $f$ is applied to the vocabulary obtained from the word-to-id mapping function Vocab($\cdot$). This process is represented by the following expression:
$\mathbf{E}_{q}=f(\mathrm{Vocab}(w_{1},w_{2},...,w_{n_{q}})),$ (1)
@@ 3.2.1 Preprocessing
Figure 3: Illustration of our path search and aggregation strategy.
@@ 3.2.2 Path Search
$\overline{\mathbf{E}}_{tri}[t][0]=\mathbf{E}_{tri}[t][0]+\mathrm{AvgPool}(\mathbf{E}_{tri}[t][0:2]),$ (2)
@@
$\begin{cases}S[t-1,j/2]=1,\;\mathrm{if}\;\epsilon_{j}>p,\\ S[t-1,j/2]=0,\;\mathrm{otherwise},\end{cases}$ (3)
@@
$\mathbf{h}=f_{t}(\overline{\mathbf{E}}_{tri}[t-1][j]||\overline{\mathbf{E}}_{tri}[t][0]),$ (4)
@@
0:  The representation of triplet pairs $\mathbf{E}_{tri}$
@@
11:           $S[\tau,j/2]=1$, $\mathbf{h}=\overline{\mathbf{E}}_{tri}[\tau][j]$
@@
3.2.3 Path Aggregation
In this section, we elaborate on the aggregation mechanism designed to use the extracted entity paths. The overarching goal is to synthesize a more refined textual representation that integrates the crux of the dialog. To achieve this, additional connections are instantiated during the backward traversal of the path, considering the inherent relationships between words. This process has been shown in Fig. 3.
Our strategy leverages an attention mechanism [43], ensuring that during the aggregation process, every node accumulates information from its antecedent nodes, thus refining its representation. For a clearer understanding, let’s consider the case of refining the $i$-th ($i\in\{0,2\}$) entity in the question. Its updated representation is computed as follows:
$\begin{cases}\mathbf{E}_{tri}^{txt}[t][i]=\overline{\mathbf{E}}_{tri}[t][i]+\sum_{j\in\mathcal{N}_{i}}\alpha_{i,j}^{t,k}\mathbf{W}_{1}\mathbf{E}_{tri}^{txt}[k][j],\\ \alpha_{i,j}^{t,k}=\mathrm{Softmax}(\gamma_{i,j}^{t,k}+1/(t-k)),\\ \gamma_{i,j}^{t,k}=\overline{\mathbf{E}}_{tri}^{T}[t][i]\mathbf{W}_{1}\mathbf{E}_{tri}^{txt}[k][j],\end{cases}$ (5)

### v4

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Information systems Multimedia information systemsccs: Computing methodologies Knowledge representation and reasoning
1. Introduction
@@
With the rapid proliferation of multimedia data, research attention has focused on vision-language tasks, proving their pivotal role in academic research (Zhang et al., 2021; Liu et al., 2023; Yin et al., 2020; Aloimonos and Fermüller, 2015; Yu et al., 2011; Xiang et al., 2024). Notable advancements have been realized in fields such as Visual Question Answering (VQA) (Liang et al., 2019; Li et al., 2021a; Yang et al., 2022; Zhang et al., 2024a, b), Visual Dialog (Guo et al., 2022; Chen et al., 2022, 2024), and Video Captioning (Seo et al., 2022; Lin et al., 2022). Central to these tasks is the aim to refine machine capabilities in understanding visual content and to convey these understandings linguistically in ways intelligible to humans (Cheng et al., 2024a, b; Lv et al., 2024a, b).
@@
An emerging domain of significance is video-grounded dialog [^1] https://video-dialog.com/., sometimes referred to as audio-visual scene-aware dialog by certain academicians (Hori et al., 2019; Le et al., 2019b; Li et al., 2021b; Zhao et al., 2023b). As illustrated in Fig. 1, this rapidly evolving domain mandates the creation of an agent adept at conducting coherent dialogs rooted in video data, which is a monumental leap toward genuine human-computer synergy. The implications of video-grounded dialogue extend widely, enhancing virtual personal assistants and creating innovative solutions for visually impaired individuals, highlighting its significant academic relevance (Li et al., 2024, 2025, 2023a). Spurred by advancements in deep learning, the domain of video-grounded dialog is witnessing considerable traction. Existing techniques in this domain can be methodically categorized into:
@@
1) Textual Information Modeling. This branch emphasizes modeling textual elements, primarily involving question inquiries and their preceding contexts (Kim et al., 2021; Le et al., 2021, 2022). For instance, Le et al. (Le et al., 2021) devised a semantic graph based on the lexical components of question-answer pairs, leading to a model proficient in forecasting reasoning pathways on the graph. This method is adept at extracting question-specific data from dialog contexts. Similarly, Kim et al. (Kim et al., 2021) employed the Gumbel-Softmax technique to identify dialog content that is most pertinent to a given question, enhancing question representation in the process.
@@
2) Visual Information Modeling. This kind of method highlights visual elements, primarily focusing on video frames and notable objects (Geng et al., 2021; Le et al., 2020; Zhang et al., 2022; Chen et al., 2023). As a case in point, Le et al. (Le et al., 2020) propounded a bidirectional model that encapsulates both spatial-to-temporal relations and their converse, aiming to mine the complex spatial and temporal cues embedded in the video. Concurrently, Geng et al. (Geng et al., 2021) advocated representing videos via spatio-temporal scene graphs, capturing salient audio-visual cues and their semantic interrelations effectively.
@@
3) Decoder Generation. This section emphasizes the formulation of answer derivations (Le and Chen, 2020; Le and Hoi, 2020; Yoon et al., 2022; Feng et al., 2023). For example, Le et al. (Le and Hoi, 2020) extended the established functionalities of the GPT-2 architecture, discarding conventional auto-regressive decoding strategy. Complementing this domain, the integration of pointer networks (Le and Chen, 2020) has amplified the generative capabilities of dialog systems.
Figure 1. An illustration of video-grounded dialog.
@@
Despite the significant progress observed in the domain of video-grounded dialog, there are still critical challenges that demand meticulous consideration: 1) Dialog History Understanding. Notable recent studies (Le et al., 2022, 2021) predominantly focus on dialog history segments that align directly with user queries. Such a perspective can inadvertently exclude broader dialog contexts, leading to the omission of significant information. The inherent lack of transparency in these methods further impedes a clear understanding of the decision-making process of the model. Hence, the imperative lies in developing a superior textual encoder capable of thoroughly analyzing dialog histories while improving model interpretability. 2) Visual Content Understanding. While several works (Geng et al., 2021) endeavor to extract visual signals through complex graph structures, the importance of user-posed questions in steering visual comprehension is frequently disregarded. This omission leads to a failure to capture question-specific insights, thereby limiting the capacity to mine meaningful visual information. Additionally, the current trend of depending solely on a singular extraction process (Zhang et al., 2022; Chen et al., 2023) may be inadequate considering the multi-faceted nature of video data. Therefore, an approach that harmonizes user queries with iterative reasoning is crucial for a nuanced comprehension of visual content.
@@
2. Related Work
2.1. Visual Dialog
@@
Visual dialog represents a sophisticated extension of the VQA framework, requiring multiple rounds of sequential interactions. The fundamental goal of this paradigm is to facilitate a seamless integration of cross-modal information derived from various inputs, culminating in the formation of a coherent textual response. Current methods within this domain may be broadly grouped into three distinct paradigms: attention mechanism-based strategies, graph network-based techniques, and approaches reliant on pre-trained models.
@@
Attention Mechanism-based Approaches. These approaches deploy attention mechanisms to pinpoint the visual regions that bear the highest relevance to the posed question. The underlying process requires an acute understanding of the immediate context of the inquiry, as well as an insight into the historical dialog trajectory. A salient example of this approach is the Recursive visual Attention (RvA) model, formulated by Niu et al. (Niu et al., 2019). Through iterative refinement of visual attention and retrospective analysis of dialog sequences, this model adeptly addresses co-reference ambiguities. Further complementing this area, Yang et al. (Yang et al., 2021) have innovated two Sequential Dialog Networks (SeqDialN). These networks artfully separate the complexity inherent in multimodal feature fusion from the inference process, thereby simplifying the design and implementation of the inference engine.
@@
Graph Network-based Techniques. Recognizing the pivotal role of seamless information integration, the academic community is increasingly embracing graph network techniques (Zhao et al., 2023a). For instance, Schwartz et al. (Schwartz et al., 2019) brought forth the Factor Graph Attention (FGA) architecture. This methodology enables the meticulous construction of graphs atop utility representations, subsequently discerning their interconnections. In a parallel vein, Guo et al. (Guo et al., 2020) introduced the Context-Aware Graph (CAG) neural structure, characterized by the dynamic interdependencies exhibited by individual nodes within the graph. Uniquely, this paradigm ensures that only nodes bearing the utmost relevance are integrated, yielding a refined, context-sensitive relational graph.
@@
Pre-trained Model-dependent Methods. The contemporary era of machine learning has been marked by the rise of pre-trained models, praised for their ability to seamlessly integrate diverse informational modalities. Anchoring on this insight, Wang et al. (Wang et al., 2020) delineated a holistic strategy that harnesses pre-trained architectures to adroitly weave together visual and dialogual elements. Mirroring this trajectory, Nguyen et al. (Nguyen et al., 2020) unveiled the Light-weight Transformer for Many Inputs (LTMI), a pioneering neural framework harnessing an agile transformer mechanism to masterfully orchestrate interactions across diverse utilities.
@@
However, the nuances and complexities of video-grounded dialog task preclude the direct application of methodologies cultivated for image-based visual dialog. Video-grounded dialog manifests longer utterance sentences, additional dialog rounds, and a more intricate context that demands sophisticated comprehension (Alamri et al., 2018). These attributes compound the challenge, rendering the direct transposition of existing visual dialog techniques impractical and inefficacious.
2.2. Video-grounded Dialog
@@
In recent years, the academic community has observed an escalating focus on video-grounded dialog systems. Noteworthy challenges, including DSTC7 (Alamri et al., 2018) and DSTC8 (Hori et al., 2020), have been instrumental in formulating benchmarks in this domain. This burgeoning field has subsequently witnessed the advent of numerous sophisticated methodologies to address its inherent complexities. For example, Hori et al. (Hori et al., 2019) pioneered an approach employing an LSTM-driven encoder-decoder framework, emphasizing multimodal attention to harmoniously integrate textual and visual cues. Expanding on this work, Le et al. (Le et al., 2019b) presented the Multimodal Transformer Network (MTN), devised specifically to encode visual content while concurrently integrating diverse data modalities. Similarly, innovative contributions like the Bi-directional Spatio-Temporal reasoning model (BiST)(Le et al., 2020) have made strides in extracting both visual and spatial cues from videos, thus elucidating the interconnectedness of text and visual sequences in multidimensional aspects. In pursuit of refining co-referential understanding, Kim et al. (Kim et al., 2021) introduced the Structured Co-reference Graph Attention model (SCGA), laying the foundation for a graph-based approach deeply rooted in multimodal co-referencing techniques. In tandem with this, Geng et al. (Geng et al., 2021) advanced the Spatio-Temporal Scene Graph Representation learning technique (STSGR)—an innovative proposition capturing the semantic crux of videos through scene graphs, all the while maintaining efficient encoding.
@@
The ascendance of pre-trained language models in myriad natural language processing domains is unmistakable. Recent efforts are directed toward fusing these models with video-grounded dialog systems. To this end, Le et al. (Le and Hoi, 2020) advanced the VGD-GPT framework, leveraging the capabilities of the well-regarded GPT-2 model, thus transforming video-grounded dialog generation into a more fluid sequence-to-sequence task. Building on this progression, Li et al. (Li et al., 2021b) incorporated a multi-task learning approach with pre-existing language models, culminating in the inception of the Response Language Model (RLM). Recognizing the hurdles of text hallucination in text generation, Yoon et al. (Yoon et al., 2022) crafted the Text HallucinAtion Mitigating (THAM) architecture, integrating Text Hallucination Regularization (THR) loss into expansive pre-trained language models. In a stride towards achieving nuanced dialog generation, Zhao et al. (Zhao et al., 2022) advocated for a Multi-Agent Reinforcement Learning (MARL) methodology, accentuating sophisticated collaborations in cross-modal analysis. With the advent of Large Video-Language Models (LVLMs), such as Video-LLaMA (Zhang et al., 2023b) and Otter (Li et al., 2023b), it is evident that this domain continues to be at the forefront of research attention.
@@
Figure 2. Illustration of our proposed framework. It consists of the textual encoder, visual encoder, and generator.
3. Methodology
As illustrated in Fig. 2, our model consists of three components: a textual encoder, a visual encoder, and a generator. The textual encoder is designed to perform interactions between questions and history contexts and output enhanced representations. Meanwhile, the visual encoder is utilized to implement progressive reasoning between questions and videos, enhancing video content comprehension. Finally, the generator combines multiple input representations to produce system answers. In the following sections, we will elaborate on each component in detail.
3.1. Problem Formulation
@@
3.2. Textual Encoder
3.2.1. Preprocessing
The word embeddings for the given question $Q_{t}$, the caption $C$, and the dialog history $H$ are obtained through the same process. For clarify, we consider the question $Q_{t}=(w_{1},w_{2},...,w_{n_{q}})$ as an example. Here, an operation $f$ is applied to the vocabulary obtained from the word-to-id mapping Vocab($\cdot$). This process is represented by the following expression:
(1) $\mathbf{E}_{q}=f(\mathrm{Vocab}(w_{1},w_{2},...,w_{n_{q}})),$
@@ 3.2.1. Preprocessing
Figure 3. Illustration of our path search and aggregation strategy.
@@ 3.2.2. Path Search
(2) $\overline{\mathbf{E}}_{tri}[t][0]=\mathbf{E}_{tri}[t][0]+\mathrm{AvgPool}(\mathbf{E}_{tri}[t][0:2]),$
@@
(3) $\begin{cases}S[t-1,j/2]=1,\;\mathrm{if}\;\epsilon_{j}>p,\\ S[t-1,j/2]=0,\;\mathrm{otherwise},\end{cases}$
@@
(4) $\mathbf{h}=f_{t}(\overline{\mathbf{E}}_{tri}[t-1][j]||\overline{\mathbf{E}}_{tri}[t][0]),$
@@
0:  The representation of triplet pairs $\mathbf{E}_{tri}$, a fully connected layer $f_{t}$
@@
11:           $S[\tau,j/2]=1$, $\mathbf{h}=f_{t}(\overline{\mathbf{E}}_{tri}[\tau][j]||\mathbf{h})$
@@
3.2.3. Path Aggregation
In this section, we elaborate on the aggregation mechanism designed to use the extracted entity paths. The overarching goal is to synthesize a more refined textual representation that integrates the crux of the dialog. To achieve this, we first reverse the constructed path during Path Search. For example, in Fig. 3, if the original path is $s_{q}\rightarrow o_{3}\rightarrow s_{1}$, after reversal, it becomes $s_{1}\rightarrow o_{3}\rightarrow s_{q}$. Then to enhance long-distance modeling capability for longer paths and enrich feature representation, we add directed edges between each node and its subsequent nodes if they do not exist. For example, in the above case, a directed edge ($s_{1}\rightarrow s_{q}$) is added from $s_{1}$ to its subsequent node $s_{q}$.
Our strategy leverages an attention mechanism (Zhang et al., 2023c; Li et al., 2022; Liu et al., 2018), ensuring that during the aggregation process, every node accumulates information from its preceding nodes based on feature similarity and distance measurement, thus refining its representation. For a clearer understanding, let’s consider the case of refining the $i$-th ($i\in\{0,2\}$) entity in the question. Its updated representation is computed as follows:
(5) $\begin{cases}\mathbf{E}_{tri}^{txt}[t][i]=\overline{\mathbf{E}}_{tri}[t][i]+\sum_{(k,j)\in\mathcal{N}_{i}}\alpha_{i,j}^{t,k}\mathbf{W}_{1}\mathbf{E}_{tri}^{txt}[k][j],\\ \alpha_{i,j}^{t,k}=\mathrm{Softmax}(\gamma_{i,j}^{t,k}+1/(t-k)),\\ \gamma_{i,j}^{t,k}=\overline{\mathbf{E}}_{tri}^{T}[t][i]\mathbf{W}_{1}\mathbf{E}_{tri}^{txt}[k][j],\end{cases}$